### v2

@@ 4.1.2 Oracle2
${\sf Oracle2}$ takes in input an integer $q\geq 1$, an instance $I$ of the SBMSm problem, a polynomial adaptive policy $\pi(t)$ that only works at round $t$ and selects all the items in $n$ steps, an integer $i\in[n]$, and picks $q$ samples $Y_{t,i}^{1},\ldots,Y_{t,i}^{q}$ of the random variable $\bm{Y}_{t,i}:=f_{t}(\{\bm{v}_{1},\ldots,\bm{v}_{i}\},\bm{\eta}_{t})-f_{t}(\{\bm{v}_{1},\ldots,\bm{v}_{i-1}\},\bm{\eta}_{t})$, that denotes the increment of $f_{t}$ achieved by $\pi(t)$ when selecting the $i$-th item $\bm{v}_{i}$ at round $t$, where $\bm{v}_{1},\ldots,\bm{v}_{n}$ are the random nodes chosen by $\pi(t)$ in order of selection. Then, ${\sf Oracle2}$ returns the empirical mean ${\sf Oracle2}(q,I,\pi(t),i):=\tilde{\Delta}_{t,i}:=\sum_{j=1}^{q}Y^{j}_{t,i}/q$.
@@
Proof.
$\leq T\cdot n\cdot 2\exp\left(-\frac{2\delta^{2}q}{\Lambda^{2}}\right)$
$=\xi.$
$\geq 1-\xi,$
and this shows property (i).
Property (ii) holds by similar arguments as in Proposition 2. ∎
@@ 4.4 Multi-Round Adaptive Policy
Theorem 1 shows that, for any fixed $\epsilon>0$ and opportune choices of $\delta$ and $\xi$, ${\sf MultiGr}_{\delta,\xi}(I)$ achieves a $(1-1/e-\epsilon)$ approximation to the optimal adaptive policy. In the following remark, we first explain why the approaches previously adopted to deal with the single-round case [26] could not be directly applied to the general multi-round framework to obtain the same approximation, thus showing that a new approach like ours was somehow necessary.
@@
Let $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},R,(f_{t})_{t\in[T]})$ be an input instance of SBMSm. The greedy multi-round policy ${\sf MultiGr}_{\delta,\xi}(I)$ is a $\left(1-1/e-\epsilon\right)$-approximate policy for $\delta:=\xi:=\frac{\lambda c\epsilon}{B(1+\Lambda)}$ and any constant $c$ such that $40<c\leq 1/$ 4, and it is polynomial in the size of $I$ and $1/\epsilon$.
Before to provide the full proof of this theorem (given in the next subsection), we would like to describe the main proof arguments necessary to get a $(1-1/e)$-approximation assuming that the probabilistic oracles are exact (i.e., assuming that Oracle1 and Oracle2 can be run with $\delta:=\xi:=0$).
Sketch of the proof (approximation factor for the case of exact oracles).
For any $k\in\llbracket 0,B\rrbracket$, let $\vec{b}(k):=(b_{1}(k),\ldots,b_{T}(k))$ be the budget vector obtained by algorithm ${\sf BudgetGr}_{\delta,\epsilon}(I)$ (see Algorithm 2) after the assignment of $k$-th unit of budget. Let ${\sf MultiGr}_{0,0}(I)$ be the greedy multi-round policy that uses exact oracles (i.e., with $\delta=\xi=0$), $\pi^{*}$ be an optimal adaptive policy, $\pi(k)$ be the greedy multi-round policy ${\sf MultiGr}_{0,0}(I)$ applied to budget vector $\vec{b}(k)$; we observe that $\pi(B)$ coincides with the policy ${\sf MultiGr}_{0,0}(I)$. Consider the following artificial adaptive policy $\hat{\pi}^{*}(k)$ that works in two phases: in the first phase, $\hat{\pi}^{*}(k)$ runs policy $\pi(k)$; in the second phase, $\hat{\pi}^{*}(k)$ restarts from round $1$ and, by observing the partial state of the previous item selections, selects further $\min\{nT-k,B\}$ items in an optimal adaptive way over all $T$ rounds.
Fix $k\in\llbracket 0,B-1\rrbracket$. By the monotonicity of the objective functions and since $\hat{\pi}^{*}(k)$ selects as many items as in the optimal adaptive policy $\pi^{*}$, the performance of $\hat{\pi}^{*}(k)$ is not worse than that of $\pi^{*}$, that is,
$\sigma(\hat{\pi}^{*}(k))\geq\sigma(\pi^{*}).$ (9)
Furthermore, by exploiting the greedy assignment of the budget chosen by ${\sf BudgetGr}_{\delta,\epsilon}(I)$ and the adaptive submodularity, one can show that $\sigma(\pi(k+1))-\sigma(\pi(k))$ (that is, the expected increment of the objective functions achieved by ${\sf MultiGr}_{0,0}(I)$, when selecting the item associated with the $(k+1)$-th unit of budget assigned by ${\sf BudgetGr}_{\delta,\epsilon}(I)$) is at least equal as the expected increment of the objective functions achieved by each item selected in the second phase of $\hat{\pi}^{*}(k)$. Thus, considering that the overall expected utility achieved in the second phase of $\hat{\pi}^{*}(k)$ is $\sigma(\hat{\pi}^{*}(k))-\sigma(\pi(k))$ and, in such phase, at most $B$ items are selected, we necessarily have that
$\sigma(\pi(k+1))-\sigma(\pi(k))\geq\frac{\sigma(\hat{\pi}^{*}(k))-\sigma(\pi(k))}{B}\geq\frac{\sigma(\pi^{*})-\sigma(\pi(k))}{B},$
where the first inequality can be formally shown by setting $\delta:=\xi:=0$ and $\pi^{\prime}(k):=\pi(k)$ in Lemma 1 (see the full proof of Theorem 1 below) and the last one follows from (9). By rearranging the above inequalities we obtain
$\sigma(\pi(k+1))\geq\left(1-\frac{1}{B}\right)\sigma(\pi(k))+\left(\frac{1}{B}\right)\sigma(\pi^{*}).$ (10)
By applying recursively (10) from $k=0$ to $k=B-1$ we obtain $\sigma(\pi(k))\geq\left(\frac{1}{B}\right)\sigma(\pi^{*})\sum_{k=0}^{B-1}\left(1-\frac{1}{B}\right)^{k}$. Thus, we have
${\sf MultiGr}_{0,0}(I)=\sigma(\pi(B))\geq\left(\frac{1}{B}\right)\sigma(\pi^{*})\sum_{k=0}^{B-1}\left(1-\frac{1}{B}\right)^{k}=\left(\frac{1}{B}\right)\sigma(\pi^{*})\left(\frac{1-\left(1-\frac{1}{B}\right)^{B}}{1-\left(1-\frac{1}{B}\right)}\right)$
$\ \ =\left(1-\left(1-\frac{1}{B}\right)^{B}\right)\sigma(\pi^{*})\geq\left(1-\frac{1}{e}\right)\sigma(\pi^{*}),$ (11)
where (11) follows from the known inequality $(1+x)^{r}<e^{rx}$, holding for any $x\geq-1$ and $r>0$, applied with $x=-1/B$ and $r=B$. By (11), we have that ${\sf MultiGr}_{0,0}(I)$ is a $(1-1/e)$-approximate policy, and this shows the claim. ∎
4.5 Full Proof of Theorem 1
@@
Fix $\delta:=\xi:=\frac{\lambda c\epsilon}{B(1+\Lambda)}$, where $c\in(0,1/4]$ is a constant that can be arbitrarily fixed. We first show the polynomial-time complexity of the algorithm. First of all, by Propositions 2 and 3 we have that both Oracle1 and Oracle2 run in polynomial time w.r.t. the size of $I$, $1/\delta$ and $1/\xi$. Then, as $1/\epsilon=O\left(\frac{nT\Lambda}{\lambda\delta}\right)=O\left(\frac{nT\Lambda}{\lambda\xi}\right)$, by the polynomial boundedness property (that polynomially relates quantities $\Lambda$ and $1/\lambda$ to the size of $I$) we conclude that Oracle1 and Oracle2 are polynomial in $1/\epsilon$ and the size of $I$. Finally, the polynomial-time complexity of the whole greedy multi-round policy follows from the fact that such policy applies $n\cdot T$ times algorithm Oracle2 to compute the budget vector (via algorithm ${\sf BudgetGr}_{\delta,\xi}(I)$) and $\sum_{t\in[T]}b_{t}=B<n\cdot T$ times the algorithm Oracle1 to select $b_{t}$ items at each round $t$ (via algorithm ${\sf SingleGr}_{\delta,\xi}(I,t,b_{t})$); the other steps of the greedy multi-round policy are less significant in terms of time-complexity.
Now, we will show the approximation guarantee of the greedy multi-round policy. Let $\pi^{*}$ denote an optimal adaptive policy for instance $I$. Let $\Delta_{t,i}$ be the expected increment of the objective function at round $t$ achieved by ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ when selecting the $i$-th item at round $t$, and let $\tilde{\Delta}_{t,i}$ be the estimation of ${\Delta}_{t,i}$ computed in algorithm ${\sf BudgetGr}_{\delta,\xi}(I)$ via Oracle2. Let ${\sf BudgetGr}^{\prime}_{\delta,\xi}(I)$ be a variant of ${\sf BudgetGr}_{\delta,\xi}(I)$ working assuming that the estimation $\tilde{\Delta}_{t,i}$ of $\Delta_{t,i}$ is correct, that is, Oracle2 (differently from Oracle1) is run with $\delta:=\xi:=0$ and then $\tilde{\Delta}_{t,i}=\Delta_{t,i}$ holds with probability $1$. For any fixed $k\in\llbracket 0,B\rrbracket$, let $\vec{b}^{\prime}(k)=(b_{1}^{\prime}(k),\ldots,b_{T}^{\prime}(k))$ be the budget vector achieved by ${\sf BudgetGr}^{\prime}_{\delta,\xi}(I)$, immediately after the assignment of $k$-th unit of budget. Let $\pi^{\prime}(k)$ denote the variant ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$ of the greedy multi-round policy that uses the budget vector $\vec{b}^{\prime}(k)=(b_{1}^{\prime}(k),\ldots,b_{T}^{\prime}(k))$ to distribute $k$ units of budget over all $T$ rounds; we observe that ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$ coincides with policy $\pi^{\prime}(B)$. Furthermore, consider the following artificial adaptive policy $\hat{\pi}^{*}(k)$ that works in two phases: in the first phase, $\hat{\pi}^{*}(k)$ runs policy $\pi^{\prime}(k)$, so that $k$ items are selected; in the second phase, $\hat{\pi}^{*}(k)$ restarts from round $1$ and, by using also the observations from the first phase, selects further $\min\{nT-k,B\}$ items in an optimal adaptive way over all $T$ rounds.
In the remainder of the proof, we will first relate the value achieved by ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$ with the optimal adaptive one achieved by $\pi^{*}$, and then we will relate ${\sf MultiGr}_{\delta,\xi}(I)$ with ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$ to obtain the desired approximation guarantee.
Fix $k\in\llbracket 0,B-1\rrbracket$. By the monotonicity of the objective functions and since $\hat{\pi}^{*}(k)$ selects as many items as in the optimal adaptive policy $\pi^{*}$, the performance of $\hat{\pi}^{*}(k)$ is not worse than that of $\pi^{*}$, that is,
$\sigma(\hat{\pi}^{*}(k))\geq\sigma(\pi^{*}).$ (12)
In the following lemma, by exploiting the adaptive submodularity, we show that $\sigma(\pi^{\prime}(k+1))-\sigma(\pi^{\prime}(k))$ (that is, the expected increment of the objective functions achieved by ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$, when selecting the item associated with the $(k+1)$-th unit of budget assigned by ${\sf BudgetGr}^{\prime}_{\delta,\epsilon}(I)$) is at least equal as $\sigma(\hat{\pi}^{*}(k))-\sigma(\pi^{\prime}(k))$ (that is, the overall expected increment of the aggregated objective function achieved in the second phase of $\sigma(\hat{\pi}^{*}(k))$) divided by $B$, up to an error $\delta+\xi\Lambda$ (due to the sampling procedure of Oracle1).

### v3

@@ 4.1.2 Oracle2
${\sf Oracle2}$ takes in input an integer $q\geq 1$, an instance $I$ of the SBMSm problem, a polynomial adaptive policy $\pi(t)$ that only works at round $t$ and selects all the items in $n$ steps, an integer $i\in[n]$, and picks $q$ samples $Y_{t,i}^{1},\ldots,Y_{t,i}^{q}$ of the random variable $\bm{Y}_{t,i}:=f_{t}(\{\bm{v}_{1},\ldots,\bm{v}_{i}\},\bm{\eta}_{t})-f_{t}(\{\bm{v}_{1},\ldots,\bm{v}_{i-1}\},\bm{\eta}_{t})$, that denotes the increment of $f_{t}$ achieved by $\pi(t)$ when selecting the $i$-th item $\bm{v}_{i}$ at round $t$, where $\bm{v}_{1},\ldots,\bm{v}_{n}$ are the random nodes chosen by $\pi(t)$ in order of selection. Then, ${\sf Oracle2}$ returns the empirical mean ${\sf Oracle2}(q,I,\pi(t),i):=\tilde{\Delta}_{t,i}:=\sum_{j=1}^{q}Y^{j}_{t,i}/q$. We have that ${\sf Oracle2}$ satisfies the following property, whose proof is deferred to Appendix C.
@@ 4.4 Multi-Round Adaptive Policy
4.5 Approximation Factor and Time-Complexity of the Policy
Theorem 2 shows that, for any fixed $\epsilon>0$ and opportune choices of $\delta$ and $\xi$, ${\sf MultiGr}_{\delta,\xi}(I)$ achieves a $1/2(1-1/e-\epsilon)$ approximation to the optimal adaptive policy and runs in polynomial-time. In the following remark, we first explain why the approaches previously adopted to deal with the single-round case [27] could not be directly applied to the general multi-round framework to obtain a constant approximation, thus showing that a new approach like ours was somehow necessary.
@@ 4.5 Approximation Factor and Time-Complexity of the Policy
Let $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},R,(f_{t})_{t\in[T]})$ be an input instance of SBMSm. The greedy multi-round policy ${\sf MultiGr}_{\delta,\xi}(I)$ is a $1/2(1-1/e-\epsilon)$-approximate policy for $\delta:=\xi:=\frac{\lambda c\epsilon}{B(4+3\Lambda)}$ and any constant $c$ such that $0<c\leq 1$, and it is polynomial in the size of $I$ and $1/\epsilon$.
Proof Sketch.
@@
We first show the polynomial-time complexity of the algorithm. First of all, by Propositions  1 and 2 we have that both Oracle1 and Oracle2 run in polynomial time w.r.t. the size of $I$, $1/\delta$ and $1/\xi$. Then, as $1/\epsilon=O\left(\frac{nT\Lambda}{\lambda\delta}\right)=O\left(\frac{nT\Lambda}{\lambda\xi}\right)$, by the polynomial boundedness property (that polynomially relates quantities $\Lambda$ and $1/\lambda$ to the size of $I$) we conclude that Oracle1 and Oracle2 are polynomial in $1/\epsilon$ and the size of $I$. Finally, the polynomial-time complexity of the whole greedy multi-round policy follows from the fact that such policy applies $n\cdot T$ times algorithm Oracle2 to compute the budget vector (via algorithm ${\sf BudgetGr}_{\delta,\xi}(I)$) and $\sum_{t\in[T]}b_{t}=B<n\cdot T$ times the algorithm Oracle1 to select $b_{t}$ items at each round $t$ (via algorithm ${\sf SingleGr}_{\delta,\xi}(I,t,b_{t})$); the other steps of the greedy multi-round policy are less significant in terms of time-complexity.
In the remainder of the proof, we focus on the analysis of the approximation factor. For simplicity, we first show that the greedy multi-round policy achieves a $1/2(1-1/e)$-approximation (that is, with $\epsilon=0$), assuming that the probabilistic oracles are exact, i.e., if the algorithms Oracle1 and Oracle2 provided in Section 4.1 can be run with $\delta:=\xi:=0$. The proof of the approximation factor for arbitrary $\delta=\xi>0$ extends the proof arguments presented below, and is given in Appendix D.
Let $\pi^{*}$ be an optimal adaptive policy for $I$, and let $OPT$ denote its expected value. Given a round $t\in[T]$ and a real number $d\geq 0$, let $\overline{OPT}_{t}(d)$ denote the optimal expected value achievable by a single-round policy at round $t$ assuming that the expected number of selected items is $d$. Instead, for an integer $b\geq 0$, let $OPT_{t}(b)$ denote the optimal expected value at round $t$ assuming that the number of selected items is equal to $b$, independently on the particular realization; we extend the above definition to any real number $d\geq 0$ by setting $OPT_{t}(d):=(d-\lfloor d\rfloor)OPT_{t}(\lceil d\rceil)+(1-(d-\lfloor d\rfloor))OPT_{t}(\lfloor d\rfloor)$, i.e., $OPT_{t}(d)$ is the expected value of an adaptive policy that optimally selects $\lceil d\rceil$ with probability $p:=d-\lfloor d\rfloor$, and optimally selects $\lfloor d\rfloor$ items with probability $1-p$. We observe that the above policy selects $d$ items in expectation, as $p\lceil d\rceil+(1-p)\lfloor d\rfloor=(d-\lfloor d\rfloor)\lceil d\rceil+(\lceil d\rceil-d)\lfloor d\rfloor=d(\lceil d\rceil-\lfloor d\rfloor)=d$ if $d\notin\mathbb{Z}_{\geq 0}$, and $p\lceil d\rceil+(1-p)\lfloor d\rfloor=pd+(1-p)d=d$ if $d\in\mathbb{Z}_{\geq 0}$; furthermore, we observe that $OPT_{t}(d)\leq\overline{OPT}_{t}(d)$ holds for any $d\geq 0$.
For any $t\in[T]$, let $d^{*}_{t}$ denote the expected number of items selected by $\pi^{*}$ at round $t$, and let $OPT_{t}$ denote the expected value of objective function $f_{t}$ under policy $\pi^{*}$. By definition of $\overline{OPT}_{t}(d^{*}_{t})$, and since states and objective functions of distinct rounds are independent, we have that
$OPT\leq\sum_{t\in[T]}\overline{OPT}_{t}(d^{*}_{t}).$ (4)
Given an integer $b\geq 0$, let $GR_{t}(b)$ denote the expected value of the greedy single-round policy ${\sf SingleGr}_{\delta,\xi}$ (see Algorithm 1) applied to round $t$, budget $b$ and $\delta=\xi=0$ (i.e., by resorting to exact oracles). Let $\vec{b}=(b_{1},\ldots,b_{T})$ denote the budget vector computed by algorithm ${\sf BudgetGr}_{\delta,\xi}$ (see Algorithm 2) with $\delta=\xi=0$, and let $GR(\vec{b})$ denote the expected value of the greedy multi-round greedy policy ${\sf MultiGr}_{\delta,\xi}$ applied to budget vector $\vec{b}$, again with $\delta=\xi=0$. We observe that
$GR(\vec{b})=\sum_{t\in[T]}GR_{t}(b_{t}).$ (5)
To show the desired approximation factor, we break down the proof into four main parts:
  1. 1.
    By resorting to (4) and Lemma 1 below, we have
    $\frac{1}{2}\cdot OPT\leq\sum_{t\in[T]}\frac{1}{2}\cdot\overline{OPT}_{t}(d^{*}_{t})\leq\sum_{t\in[T]}OPT_{t}(d^{*}_{t}).$ (6)
    We stress that the proof of this step uses the strong submodularity property (in particular, it is used in the proof of Lemma 1 below).
  2. 2.
    By derandomizing vector $\vec{d}^{*}:=(d^{*}_{1},\ldots,d^{*}_{T})$, we have that there exists a budget vector $\vec{b^{*}}=(b_{1}^{*},\ldots,b_{T}^{*})$ of non-negative integers with $\sum_{t\in[T]}b^{*}_{t}=B$ such that
    $\sum_{t\in[T]}OPT_{t}(d^{*}_{t})\leq\sum_{t\in[T]}OPT_{t}(b^{*}_{t})$ (7)
    (see Lemma 2 below).
  3. 3.
    Then, by exploiting the $(1-1/e)$-approximation of the standard adaptive greedy algorithm [27] (that resorts to the adaptive submodularity property), we have
    $\sum_{t\in[T]}\left(1-\frac{1}{e}\right)OPT_{t}(b_{t}^{*})\leq\sum_{t\in[T]}GR_{t}(b^{*}_{t})=GR(\vec{b}^{*})$ (8)
    (see Lemma 3 below).
  4. 4.
    Finally, by exploiting the greedy assignment of ${\sf BudgetGr}_{0,0}$ and the adaptive submodularity, we have
    $GR(\vec{b}^{*})\leq GR(\vec{b}).$ (9)
    (see Lemma 4 below).
By combining the above inequalities we obtain
$\frac{1}{2}\left(1-\frac{1}{e}\right)OPT$ $\leq\left(1-\frac{1}{e}\right)\sum_{t\in[T]}OPT_{t}(d^{*}_{t})\leq GR(\vec{b}^{*})\leq GR(\vec{b}),$ (10)
and this shows the claim. ∎
4.6 Technical Lemmas
Below are the lemmas cited in the proof sketch of Theorem 2.